OpenStudy (anonymous):

8t + 8 t^2 - 1 ----------- * ------------ 2t^2 + t - 1 t^2 - 2t + 1

OpenStudy (owlcoffee):

The multiplication of two fractions is the product of denominators and numerators: \[\frac{ a }{ b }\times \frac{ x }{ y } \iff \frac{ ax }{ by }\]

OpenStudy (anonymous):

so it would be?